Guerlane Fragrance Corporation has a perfume division, Essense, and a cologne division, Karisma. The following information relates to last year's operations at Karisma: Net operating income = $10,000

Profit Margin = 5%

Investment Turnover = 2.5

Minimum required rate of return = 15%

Calculate Total Assets.

Calculate Residual income and is Karisma division meeting Guerlane Fragrance Corporation's required rate of return.
Answer #1

Profit Margin = Net operating income/Sales

5% = 10,000/Sales

Sales = 10,000/5%

= $200,000

Investment turnover = Sales/Total assets

2.5 = 200,000/Total assets

Total assets = 200,000/2.5

= $80,000

Residual income = Net operating income - (Average operating assets x Required rate of return)

= 10,000 - (80,000 x 15%)

= 10,000 - 12,000

= - $2,000

Since residual income of Karisma division is negative, hence Karisma division is not meeting Guerlane Fragrance Corporation's required rate of return.

Kindly comment if you need further assistance. Thanks